OpenStudy (anonymous):

What percent of 250 is 70? (1 point) 3.6% 18% 28% 175%

jimthompson5910 (jim_thompson5910):

(x/100)*250 = 70 x/100 = 70/250 x = (70/250)*(100) x = ???

OpenStudy (anonymous):

The answer is C. (28%).

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

No problem.